OpenStudy (anonymous):

Write an equation of a line that has a slope of 1/2 and passes through (2,1) .

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] where m is the gradient/slope of the line. and \[(x _{1},y _{1})\] is the point where the line passes.

OpenStudy (anonymous):

\[y-1=\frac{ 1 }{ 2 }(x-2)\]

OpenStudy (anonymous):

expand or multiply both sides by 2. Depending whether you want it in general form ax+by+c=0 or in gradient intercept form: y=mx+b

OpenStudy (anonymous):

thanks !